OpenStudy (anonymous):

Use a table to find the percentage of the area under the standard normal curve between the two values. Round your answer to the nearest tenth. z = 0 and z = 1.89

OpenStudy (anonymous):

What kind of table? The probability Z table?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Well using this: http://4.bp.blogspot.com/_5u1UHojRiJk/TEh9BHxxPUI/AAAAAAAAAIQ/DafeQNMYFoE/s1600/ztable.gif 0 on top, and ~1.89 on side, so 0.9713 sound about right?

OpenStudy (anonymous):

No using this one 2nd page

OpenStudy (anonymous):

Hmm, alright, lets see... z=0 is 0.5, z= 1.89 is 0.9706 So if you want the area between them, 0.9706-0.5=0.4706, so 47.06%?

OpenStudy (anonymous):

That's right thank you so much!! :) Can you help me with one more? Use a z-table to find the percentage of the area under the standard normal curve between the two values. Round your answer to the nearest hundredth. z = –0.32 and z = 1.94

OpenStudy (anonymous):

z=-.32 is 0.3745 and z=1.94 is 0.9738 So.. we want what is between them... 0.9735-.5 = 0.4735 0.5-0.3745=0.1225 ^^This gets the distance from the center point 0.5 on either side 0.4735+0.1225=0.596 or 59.6% if i did the math right?